The floor area of a rectangular classroom measures 60 meter squared. The floor is to be covered with square tiles measuring 30 cm by 30 cm. How many tiles will it take to cover the floor? Assume that only the whole tiles are used. Give a short, written description of how you solved the problem
Found 2 solutions by mananth, ewatrrr:
Answer by mananth(16946)

area of class room = 60 m^2
..
tile size = 30 cm by 30cm
The tiles are in cm
convert them to meters
1meter = 100 cm
30 cm = 0.3 m
Area of tile = 0.3*0.3= 0.09m^2
..
The area of the classroom divided by the area of the tile will give us the number of tiles required .
60/0.09
=667 tiles will be required.
...
m.ananth@hotmail.ca

Answer by ewatrrr(24785)

Hi,
.
question states
floor area = 60m^2
.
*Note: 30cm = .3m
Area of each tile in m^2:
.3^2 = .09m^2
.
60m^2 / .09 m^2 = 666.67
.
would take 667 whole tiles.